OpenStudy (anonymous):

Determine if triangle DEF with coordinates D (2, 1), E (3, 5), and F (6, 2) is an equilateral triangle. Use evidence to support your claim. If it is not an equilateral triangle, what changes can be made to make it equilateral? Be specific.

OpenStudy (anonymous):

@djcool31

OpenStudy (anonymous):

@undeadknight26

OpenStudy (jadeishere):

1st step, get to a graphing site, or use graphing paper to make a coordinate plane

OpenStudy (jadeishere):

2nd step, graph the coordinates

OpenStudy (anonymous):

ok

OpenStudy (jadeishere):

hold on, let me graph it real quick ;D

OpenStudy (jadeishere):

3rd step, look at all of the angles, and sides of the triangle, measure the distance between each corner

OpenStudy (anonymous):

how do you do that

OpenStudy (jadeishere):

well, an easy way is just to get a ruler and put it up to the screen, or, if you have a good graphing site or something to use, you can measure the angles. I recommend Geogebra.org

OpenStudy (anonymous):

how can I do it with math or showing work

OpenStudy (jadeishere):

This might help, it's the graph I made, you can look at the angles, and see they're not exactly congruent

OpenStudy (anonymous):

I thought that an equilateral just had to have the same lengths

OpenStudy (jadeishere):

It does, but the angles have to do with the length itself. Angles are what makes triangles, well.. triangles

OpenStudy (anonymous):

true so do you know how to do that without a graphing thing

OpenStudy (jadeishere):

do you have a protractor?

OpenStudy (anonymous):

no I need to do it for a flvs question and I don't remaber how to do any of this stuff

OpenStudy (jadeishere):

if you want to measure the angle of a triangle, I think you need a protractor.. let's see if we can make something up.. First of all, measure the lengths of the sides, tell me what you get, please

OpenStudy (jadeishere):

So it's like this a = 4.24 b = 4.12 c = 4.12 So this is an Isosceles triangle

OpenStudy (anonymous):

is there a formula to do that with

OpenStudy (jadeishere):

There probably is, I forgot it, Hold on, i'll look through my notes

OpenStudy (anonymous):

ok

OpenStudy (jadeishere):

cosA = (b^2 + c^2 − a^2) over 2bc cosB = (c^2 + a^2 − b^2) over 2ca cosC = (a^2 + b^2 − c^2) over 2ab So, that's the cosine formula to get angles When I say 'over 2bc, 2ca or 2ab' I mean it's b^2 + c^2 = a^2 --------------- 2bc

OpenStudy (anonymous):

i was taught that you use distends formal i know it has a square root in it that is it

OpenStudy (jadeishere):

Use this formula if you want to get the angle. it's easy. cosA =b2 + c2 − a2 ----------- 2bc So cosA = 4.12^2 + 4.12^2 − 4.24^2 ------------------------ 2 * 4.12 * 4.12 cosA = 16.97 + 16.97 - 17.98 -------------------- 33.95 cosA= 15.96 ______ 33.95 cos A = 0.47010 degrees if i did that right for A, the angle is 0.47010 I DID round for some of the decimals

OpenStudy (anonymous):

Ok

OpenStudy (jadeishere):

Yep, so, in answer to the actual question, no it is not an equilateral because B and C have the same length, but A is not equal, so it IS an isosceles. The way you could change it to be an equilateral is if you shortened the size of A by 0.12

OpenStudy (anonymous):

ok thanks

OpenStudy (jadeishere):

By the way, D = A, E = B and F = C